# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ 2.2 Incorporating survey weights into propensity score weighting
In this equation, the weights are given by
@@
Details of proof are included in Web Appendix A.1. It is important to note that the above derivation primarily applies to retrospective studies, where exposure occurs before survey sampling begins. However, this approach can also be generalized to prospective studies by defining the sampling weights as $p(S_{i}=1\mid\mathbf{X}_{i})$ for individual $i$. In such cases, one can factorize $p(Z_{i}=1,S_{i}=1\mid\mathbf{X}_{i})$ differently as $p(S_{i}=1\mid\mathbf{X}_{i})p(Z_{i}=1\mid S_{i}=1,\mathbf{X}_{i})$, where the second term represents the in-sample propensity score.[14] Further details on the corresponding estimators under prospective settings are extended and discussed in Web Appendix A.6.
@@
The specification of $h(\mathbf{X})$ can adopt various forms depending on specific statistical or scientific considerations. Some commonly used forms of balancing weights and their associated target populations, tilting functions, and estimands are presented in Table 1. In this table, PATE refers to the Population Average Treatment Effect, defined as $\tau_{\text{PATE}}=E[Y(1)-Y(0)]$, with $g(\mathbf{X})=f(\mathbf{X})$ and $h(\mathbf{X})=1$. PATT denotes the Population Average Treatment Effect for the Treated, expressed as $\tau_{\text{PATT}}=E[Y(1)-Y(0)\mid Z=1]$, where $g(\mathbf{X})=f(\mathbf{X}\mid Z=1)$ and $h(\mathbf{X})=e_{\text{sp}}(\mathbf{X})$. Complimentary to PATT, PATC represents the Population Average Treatment Effect for the Control, given by $\tau_{\text{PATC}}=E[Y(1)-Y(0)\mid Z=0]$, with $g(\mathbf{X})=f(\mathbf{X}\mid Z=0)$ and $h(\mathbf{X})=1-e_{\text{sp}}(\mathbf{X})$. The overlap weights defined by Li et al. [9] correspond to the Average Treatment Effect among the Overlap Population (PATO) under the tilting function of the average $h(\mathbf{X})=e_{\text{sp}}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)$, such that greater emphasis is given to units with propensity scores close to 0.5 and smaller weights were assigned to those with extreme propensity scores with more deterministic treatment assignment. In the absence of survey sampling, the choice of the overlap tilting function attains the minimum asymptotic variance among all choices of the titlting functions under the homoscedasticity assumption. [9] Under the survey observational data setting, we show that this minimum asymptotic variance property continues to hold, along with the exact balance properties. We summarize the properties below and details of the proof can be found in Web Appendix A.2.
@@
When propensity scores are estimated using a survey-weighted logistic regression under weighted maximum likelihood estimation, where $\widehat{e}_{\text{sp}}(\mathbf{X}_{i})=\operatorname{logit}^{-1}(\widehat{\boldsymbol{\beta}}_{0}+\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}_{sw}^{\top})$, the application of overlap weights ensures an exact balance in the population means of any given covariate across treatment and control groups. That is,
$\frac{\sum_{i=1}^{n}X_{ik}Z_{i}(1-\widehat{e}_{\text{sp}}(\mathbf{X}_{i}))/p_{1}(\mathbf{X}_{i})}{\sum_{i=1}^{n}Z_{i}(1-\widehat{e}_{\text{sp}}(\mathbf{X}_{i}))/p_{1}(\mathbf{X}_{i})}=\frac{\sum_{i=1}^{n}X_{ik}(1-Z_{i})\widehat{e}_{\text{sp}}(\mathbf{X}_{i})/p_{0}(\mathbf{X}_{i})}{\sum_{i=1}^{n}(1-Z_{i})\widehat{e}_{\text{sp}}(\mathbf{X}_{i})/p_{0}(\mathbf{X}_{i})},\quad\text{for}\quad k=1,2,\ldots,K$
Proposition 2.
The function $h(\mathbf{X})\propto e_{\text{sp}}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)$ minimizes the asymptotic variance of the weighting estimator $\widehat{\tau}_{h}$ across all possible $h$ functions, under the assumption of homoscedasticity for both the original outcome $Y(z)$ and the pseudo-outcome $Y(z)/p_{z}(\mathbf{X})$ across treatment groups ($z=1$ and $z=0$). As the sample size $n$ approaches infinity:
$n\min_{h}\left\{E_{\mathbf{X}}\text{Var}\left[\widehat{\tau}_{h}\mid\mathbf{X},S=1\right]\right\}\to\frac{\sqrt{vc}}{C_{hs}^{2}}\int e_{\text{sp}}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)\frac{1}{p(S=1\mid\mathbf{X})}f(\mathbf{X}\mid S=1)\mu(d\mathbf{X}),$
where $v=\text{Var}\left(Y(z)/p_{z}(\mathbf{X})\mid\mathbf{X}\right)$, $c=\text{Var}\left(Y(z)\mid\mathbf{X}\right)$, and $C_{hs}=\int(h(\mathbf{X})/p(S=1\mid\mathbf{X}))f(\mathbf{X}\mid S=1)\mu(d\mathbf{X})$.
Proposition 2 leads to optimal efficiency in estimating $\widehat{\tau}_{h}$, minimizing the variance among all forms of balancing weights. While the homoscedasticity condition may not always hold in practice, our simulations demonstrate that using overlap weights consistently achieves smaller variances in finite survey samples. Furthermore, previous empirical work on nonsurvey-weighted outcomes has shown that overlap weights improve efficiency compared to inverse probability of treatment weighting (IPTW), regardless of the homoscedasticity condition.[19, 20, 21] Under these two propositions, overlap weights can still mitigate the covariate balancing and extreme propensity score issues under the survey sampling setup. In addition, for truncated combined, the tilting function $h(\mathbf{X})=1\{\alpha<e_{\text{sp}}(\mathbf{X})<1-\alpha\}$ can be used to exclude individuals with extreme propensity scores, which facilitates more reliable estimations by focusing on a population with sufficient overlap.[22] The parameter $\alpha$, typically chosen within the range $(0,0.1)$, indicates the truncation threshold for propensity scores.
Table 1: Survey-weighted propensity score weights for different forms of balancing weights, target population, tilting function, and estimand.
[…]
@@ 2.3 Asymptotic normality and the sandwich variance estimator
Under the framework of balancing weights, this section establishes the asymptotic normality of the propensity score weighting estimator and derives the robust sandwich variance estimator in the presence of survey weights. One viable approach for inference is the bootstrap method, which could ascertain the variance of $\widehat{\tau}_{h}$ across various target populations. However, bootstraps can involve a substantial amount of computation, especially when the sample size gets larger. In addition, bootstrap survey data often entails additional complications, see, for example, Ackerman et al.[23] Leveraging the asymptotic normality property of the survey-weighted estimator, we provide a closed-form variance estimator via the empirical sandwich method, which serves as a feasible solution to circumvent the need for extensive computation.[24] Detailed proof can be found in Web Appendix A.3.
Assuming the propensity scores are estimated using a survey-weighted logistic regression model and in the form of $e_{\text{sp}}\left(\mathbf{X}_{i};\boldsymbol{\beta}\right)=1/\left(1+\exp\left(-\mathbf{X}_{i}^{\top}\boldsymbol{\beta}\right)\right)$. We define $\tau_{1}=\displaystyle\frac{E\left[h(\mathbf{X})\mu_{1}(\mathbf{X})\right]}{E\left[h(\mathbf{X})\right]}$and $\tau_{0}=\displaystyle\frac{E\left[h(\mathbf{X})\mu_{0}(\mathbf{X})\right]}{E\left[h(\mathbf{X})\right]}$, where $\mu_{1}(\mathbf{X})=E\left[Y(1)\mid\mathbf{X}\right]$ and $\mu_{0}(\mathbf{X})=E\left[Y(0)\mid\mathbf{X}\right]$, as the treatment-specific average potential outcomes in the target population, and the treatment effect estimator of target population can be written as $\widehat{\tau}_{h}=\widehat{\tau}_{1}-\widehat{\tau}_{0}$. The consistent sandwich-type variance estimator for $\widehat{\tau}_{h}$ is given by:
@@
where $\widehat{\nu}=n^{-1}\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})$ represents the weighted adjustment factor that accounts for survey weights $1/p_{z}(\mathbf{X}_{i})$. The influence function $\widehat{I}_{i}$ here, characterizing the contribution of the $i$-th sampled individual to the overall population-level treatment effect estimator, is defined as:
@@
The sandwich variance estimator accounts for the uncertainty in the estimation of propensity scores and also adjusts for the known survey weights.
@@ 3 Improving the weighting estimator through outcome regression
To improve the weighting estimators, augmented weighting estimators have proven useful for causal inference and missing data analysis.[7, 25, 26] These estimators are specifically designed to leverage additional information from outcome modeling, and for inverse probability weighting (IPW), can provide consistent estimation of causal effects even if only one of the models—treatment assignment mechanism or the outcome model——is correctly specified, but not necessarily both. In the absence of survey sampling, Mao et al.[27] investigated these augmentation methods to estimate different target estimands in observational studies under the balancing weight framework.
In this section, we first develop the augmented weighting (AW) estimating equations based on the balancing weights in Section 3.1 in survey observational data. We then introduce three versions of augmented estimators adapted to in Sections 3.2 to 3.4: the moment estimator, the clever covariate estimator, and the weighted regression estimator. Each augmentation weighting method is discussed with a focus on how survey weights are integrated at various stages, clarifying their applications in the survey setting. Section 3.5 presents the closed-form sandwich variance estimation for the augmented estimators for computationally efficient inference. This section primarily focuses on the development of estimators under the retrospective case, consistent with our previous notation, while modifications required for prospective designs are outlined in the Web Appendix A.6.
@@ 3.1 Augmented weighting estimating equations for survey observational data
Following the framework outlined in §13.5 of Tsiatis [28] for defining the class of influence functions for regular asymptotically linear estimators, we can extend the asymptotic properties of augmented estimators[27] that incorporate survey weights. In Web Appendix A.4, we prove the proposition below that guarantees the local efficiency of the augmented estimator after incorporating the survey weights.
@@
If the survey weights and the propensity scores model are known, the class of observed sample data influence functions for regular asymptotically linear estimators of population estimand is,
$\tau_{h}^{\text{AW}}=\left\{\frac{{h(e_{\text{sp}}(\mathbf{X}))}/{P(S=1\mid\mathbf{X})}}{\Psi}\left[\displaystyle\frac{ZY}{e_{\text{fp}}(\mathbf{X})}-\displaystyle\frac{(1-Z)Y}{1-e_{\text{fp}}(\mathbf{X})}\right]-\tau_{h}\right\}+\{Z-e_{\text{fp}}(\mathbf{X})\}\varphi(\mathbf{X}),$
for any function $\varphi(\mathbf{X})$.
@@
Here $\Psi=E\left[\left.{h(e_{\text{sp}}(\mathbf{X}))/{P(S=1\mid\mathbf{X})}}\right|S=1\right]$ is the normalization term and we use $h(e_{\text{sp}}(\mathbf{X}))$ to explicitly highlight that the tilting function $h(\mathbf{X})$ can depend on the population propensity score $e_{\text{sp}}(\mathbf{X})$, as illustrated in Table 1. Based on this proposition, one can follow the argument in Tsiatis and show that, among all estimators whose influence functions fall within this class, the augmented estimator with
$\varphi(\mathbf{X})$ $=-\frac{h(e_{\text{sp}}(\mathbf{X}))/p(S=1\mid\mathbf{X})}{\Psi}\left[\frac{E[Y\mid Z=1,\mathbf{X},S=1]}{e_{\text{fp}}(\mathbf{X})}+\frac{E[Y\mid Z=0,\mathbf{X},S=1]}{1-e_{\text{fp}}(\mathbf{X})}\right]$ (9)
@@
where $E\left[Y\mid\mathbf{X},Z=1,S=1\right]$ and $E\left[Y\mid\mathbf{X},Z=0,S=1\right]$ are modeled by correctly specified outcome models, achieves the smallest asymptotic variance.[28] This theoretical guarantee allows us to obtain the augmented weighting estimating equation tailored to survey observational data:
$\frac{1}{\Psi}\frac{1}{n}\sum_{i=1}^{n}\Bigg{\{}\frac{h(\mathbf{X}_{i})}{P(S_{i}=1\mid\mathbf{X}_{i})}\big{(}m_{1}(\mathbf{X}_{i})-m_{0}(\mathbf{X}_{i})\big{)}$ $+\frac{Z_{i}h(\mathbf{X}_{i})}{e_{\text{sp}}(\mathbf{X}_{i})p_{1}(\mathbf{X}_{i})}\big{(}Y_{i}-m_{1}(\mathbf{X}_{i})\big{)}$
$-\frac{(1-Z_{i})h(\mathbf{X}_{i})}{(1-e_{\text{sp}}(\mathbf{X}_{i}))p_{0}(\mathbf{X}_{i})}\big{(}Y_{i}-m_{0}(\mathbf{X}_{i})\big{)}-\tau_{h}^{\text{AW}}\Bigg{\}}=0,$ (10)
where $m_{1}(\mathbf{X})=E\left[Y(1)\mid\mathbf{X}\right]=E\left[Y\mid\mathbf{X},Z=1,S=1\right]$ represents the conditional expectation of the outcome given covariates for the treatment group in the survey sample. The second term in Equation (3.1) relies on the weak exchangeability assumption for both $Z$ and $S$, as discussed in Section 2.1. Similarly, $m_{0}(\mathbf{X})=E\left[Y(0)\mid\mathbf{X}\right]=E\left[Y\mid\mathbf{X},Z=0,S=1\right]$ is the conditional expectation for the control group. This formulation enables the estimation of the treatment effect between potential outcomes in the population based on observed data under the assumption of weak exchangeability. By plugging in $\Psi$ in Equation (3) as $\Psi=E\left[Z\omega_{1,p}(\mathbf{X})\mid S=1\right]=E\left[(1-Z_{i})\omega_{0,p}(\mathbf{X})\mid S=1\right]=E\left[h(\mathbf{X})/{p(S=1\mid\mathbf{X})}\mid S=1\right]$, the general solution to this augmented weighting estimating equation becomes,
$\widehat{\tau}_{h}^{\text{AW}}$ $=\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})\left(\widehat{m}_{1}(\mathbf{X}_{i})-\widehat{m}_{0}(\mathbf{X}_{i})\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})}$ (11)
$\quad+\frac{\sum_{i=1}^{n}Z_{i}\omega_{1,p}(\mathbf{X}_{i})\left(Y_{i}-\widehat{m}_{1}(\mathbf{X}_{i})\right)}{\sum_{i=1}^{n}Z_{i}\omega_{1,p}(\mathbf{X}_{i})}-\frac{\sum_{i=1}^{n}(1-Z_{i})\omega_{0,p}(\mathbf{X}_{i})\left(Y_{i}-\widehat{m}_{0}(\mathbf{X}_{i})\right)}{\sum_{i=1}^{n}(1-Z_{i})\omega_{0,p}(\mathbf{X}_{i})}$ (12)
where $(\omega_{1,p}(\mathbf{X}_{i}),\omega_{0,p}(\mathbf{X}_{i}))$ are balancing weights pair for survey observational data, as defined in equation (4) in Section 2.2. In retrospective studies, when the marginalized survey weights $P(S_{i}=1\mid\mathbf{X}_{i})$ are not directly accessible, they can be approximated using $r_{z}(\mathbf{X}_{i})p_{z}(\mathbf{X}_{i})=P(S_{i}=1\mid\mathbf{X}_{i})$ for each individual with $z_{i}=1$ or 00.
This estimator is a locally efficient estimator for $\tau_{h}$, and the details of the proof can be found in Web Appendix A.4. The consistency of $\widehat{\tau}_{h}^{\text{AW}}$ is contingent on the inclusion of the population-level propensity score $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$ in the tilting function $h(\mathbf{X})$ in the first term of Equation (11). If the population-level propensity score model is mis-specified with probability limit defined as $e_{\text{sp}}(\mathbf{X};\widetilde{\boldsymbol{\beta}})$, the estimator will inevitably be biased for the original weighted average treatment effect estimand but still consistent for $\widetilde{\tau}_{h}={E\left[h\left(e_{\text{sp}}(\mathbf{X};\widetilde{\boldsymbol{\beta}})\right)\tau(\mathbf{X})\right]}/{E\left[h(e_{\text{sp}}(\mathbf{X};\widetilde{\boldsymbol{\beta}}))\right]}$, which represents the average treatment effect in a shifted target population determinded by the shifted tilting function $h(e_{\text{sp}}(\mathbf{X};\widetilde{\boldsymbol{\beta}}))$. Generalizing from Mao et al.,[27] we have the following proposition (detailed derivations in Web Appendix A.4) .

### v2

@@ 2.2 Incorporating survey weights into propensity score weighting with balancing weights
with weights given by
@@
The derivation is included in Web Appendix A.1. Although this expression is specialized to retrospective designs, it can be modified to address prospective designs by defining the sampling weights as $p(S_{i}=1\mid\mathbf{X}_{i})$ for individual $i$. In such cases, one can factorize $p(Z_{i}=1,S_{i}=1\mid\mathbf{X}_{i})$ differently as $p(S_{i}=1\mid\mathbf{X}_{i})p(Z_{i}=1\mid S_{i}=1,\mathbf{X}_{i})$, where the second term represents the sample-level propensity score.[14] Further details on the corresponding estimators under prospective designs are discussed in Web Appendix A.6.
The specification of $h(\mathbf{X})$ can adopt various forms depending on specific statistical or scientific considerations. Some commonly used forms of balancing weights and their associated target populations, tilting functions, and estimands are presented in Table 1. For example, we obtain PATE defined as $\tau_{\text{PATE}}=E[Y(1)-Y(0)]$, when $g(\mathbf{X})=f(\mathbf{X})$ and $h(\mathbf{X})=1$. We obtain PATT expressed as $\tau_{\text{PATT}}=E[Y(1)-Y(0)\mid Z=1]$, when $g(\mathbf{X})=f(\mathbf{X}\mid Z=1)$ and $h(\mathbf{X})=e_{\text{sp}}(\mathbf{X})$. Complementary to PATT, PATC represents the Population Average Treatment Effect for the Control, given by $\tau_{\text{PATC}}=E[Y(1)-Y(0)\mid Z=0]$, with $g(\mathbf{X})=f(\mathbf{X}\mid Z=0)$ and $h(\mathbf{X})=1-e_{\text{sp}}(\mathbf{X})$. Importantly, these different tilting functions target subpopulations within the original external population from which the survey sample is drawn and hence the actual target population can differ from that original survey population. In our context, the survey weights are used to inform the original external population from which the survey sample is drawn, whereas $h(\mathbf{X})$ determines which subpopulation receives emphasis in the causal effect estimation. Even when the external population is well defined through a probability sampling design, causal identification from observational data still hinges on adequate overlap in treatment assignment within that population. The overlap and truncated combined estimands below formalize this principle. Specifically, for the truncated combined approach, the tilting function $h(\mathbf{X})=1\{\alpha<e_{\text{sp}}(\mathbf{X})<1-\alpha\}$ can be used to exclude individuals with extreme propensity scores and to target a population with sufficient overlap.[16, 17] This defines a “region of common support” within the original external population where treatment effects are identifiable without extrapolation, while preventing the multiplicative impact due to extreme survey weights $1/p_{z}(\mathbf{X})$ and extreme propensity score weights $1/e_{sp}(\mathbf{X})$ or $1/(1-e_{sp}(\mathbf{X}))$ that would otherwise result in bias and inflate variance. The parameter $\alpha$, typically chosen within the range $(0,0.1)$, indicates the truncation threshold for propensity scores.
@@
The overlap weights defined by Li et al. [9] correspond to the Average Treatment Effect among the Overlap Population (PATO) under the tilting function of the average $h(\mathbf{X})=e_{\text{sp}}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)$, such that greater emphasis is given to units with propensity scores close to 0.5 and smaller weights are assigned to those with extreme propensity scores with more deterministic treatment assignment. In survey observational studies, this overlap population is defined relative to the survey’s original external population and still includes individuals whose treatment decisions are genuinely uncertain—the equipoise region where both treatments are plausible options and treatment effect evidence is most valuable for informing policy or clinical decisions. With additional regularity assumptions, we show that the original minimum asymptotic variance property of overlap weights continues to hold, along with the exact balance property under survey-weighted logistic propensity score modeling. We summarize the properties below and the proof can be found in Web Appendix A.2.
Table 1: A summary of different forms of balancing weights, target population, tilting function, and estimand for survey observational studies.
[…]
@@
When the propensity scores are estimated using a survey-weighted logistic regression under weighted pseudo maximum likelihood estimation, where $\widehat{e}_{\text{sp}}(\mathbf{X}_{i})=\operatorname{logit}^{-1}(\widehat{\boldsymbol{\beta}}_{0}+\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}_{sw}^{\top})$, the application of overlap weights ensures exact mean balance of any given covariate between groups in the target population. That is,
$\frac{\sum_{i=1}^{n}X_{ik}Z_{i}(1-\widehat{e}_{\text{sp}}(\mathbf{X}_{i}))/p_{1}(\mathbf{X}_{i})}{\sum_{i=1}^{n}Z_{i}(1-\widehat{e}_{\text{sp}}(\mathbf{X}_{i}))/p_{1}(\mathbf{X}_{i})}=\frac{\sum_{i=1}^{n}X_{ik}(1-Z_{i})\widehat{e}_{\text{sp}}(\mathbf{X}_{i})/p_{0}(\mathbf{X}_{i})}{\sum_{i=1}^{n}(1-Z_{i})\widehat{e}_{\text{sp}}(\mathbf{X}_{i})/p_{0}(\mathbf{X}_{i})},\quad\text{for}\quad k=1,2,\ldots,K.$
In the absence of survey sampling or under simple random sampling, overlap weights also possess the optimal efficiency property, as proved in Li et al.[9] Under a retrospective survey sampling design, in Web Appendix A.2 (Proposition S1), we prove that when the standard homoskedasticity assumption holds, that is, $\text{Var}(Y(z)\mid\mathbf{X})=\sigma^{2}$, the theoretically variance-minimizing tilting function is in fact $h^{*}(\mathbf{X})\propto\left(p_{1}(\mathbf{X})p_{0}(\mathbf{X})/p(\mathbf{X})\right)e_{\text{sp}}(\mathbf{X})(1-e_{\text{sp}}(\mathbf{X}))$, which depends on the treatment-specific sampling probabilities $p_{1}(\mathbf{X})$ and $p_{0}(\mathbf{X})$. While this tilting function is theoretically appealing, it is less interpretable in practical applications. On the other hand, the original tilting function that generates the standard overlap weights, $h(\mathbf{X})\propto e_{\text{sp}}(\mathbf{X})(1-e_{\text{sp}}(\mathbf{X}))$, attains the minimum variance property under two conditions: (i) simple random sampling where $p_{1}(\mathbf{X})=p_{0}(\mathbf{X})=p$ is constant, with the standard homoskedasticity assumption, or (ii) covariate-dependent sampling with a modified homoskedasticity condition $\text{Var}(Y(z)/\sqrt{p_{z}(\mathbf{X})}\mid\mathbf{X})=\tilde{\sigma}^{2}$. Due to interpretability, we focus on the standard overlap weights throughout. While the modified homoskedasticity condition in (ii) is primarily for theoretical purposes and may not hold in practice, our simulations in Section 4 demonstrate that overlap weights can still achieve smaller variances than inverse probability weights in finite samples.
@@ 2.3 Asymptotic normality and the sandwich variance estimator
One viable approach for statistical inference is the bootstrap method, which could ascertain the variance of $\widehat{\tau}_{h}$ across various target populations. However, bootstraps can involve a substantial amount of computation, especially when the sample size gets larger. In addition, bootstrap survey data often entails additional complications, see, for example, Ackerman et al.[18] We provide a closed-form variance estimator via the empirical sandwich method, which serves as a feasible solution to circumvent the need for extensive computation.[19] Detailed derivations can be found in Web Appendix A.3.
Assuming the propensity scores are estimated using a survey-weighted logistic regression model and in the form of $e_{\text{sp}}\left(\mathbf{X}_{i};\boldsymbol{\beta}\right)=1/\left(1+\exp\left(-\mathbf{X}_{i}^{\top}\boldsymbol{\beta}\right)\right)$. We define $\tau_{1}={E\left[h(\mathbf{X})\mu_{1}(\mathbf{X})\right]}/{E\left[h(\mathbf{X})\right]}$ and $\tau_{0}={E\left[h(\mathbf{X})\mu_{0}(\mathbf{X})\right]}/{E\left[h(\mathbf{X})\right]}$, where $\mu_{1}(\mathbf{X})=E\left[Y(1)\mid\mathbf{X}\right]$ and $\mu_{0}(\mathbf{X})=E\left[Y(0)\mid\mathbf{X}\right]$, as the treatment-specific average potential outcomes in the target population, and the treatment effect estimator of target population can be written as $\widehat{\tau}_{h}=\widehat{\tau}_{1}-\widehat{\tau}_{0}$. The consistent sandwich-type variance estimator for $\widehat{\tau}_{h}$ is given by:
@@
where $\widehat{\nu}=n^{-1}\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})$. Here, the influence function $\widehat{I}_{i}$ characterizing the contribution of the $i$-th sampled individual to the effect estimator is defined as:
@@
This sandwich variance estimator extends the one in Li et al.[20] by accounting for the known survey weights and the uncertainty in the estimation of survey-weighted propensity scores.
@@ 3 Improving the weighting estimator through outcome regression
In this section, we first develop the augmented weighting (AW) estimating equations based on the balancing weights in Section 3.1 in survey observational data. We then introduce three versions of augmented estimators in Sections 3.2 to 3.4: the one-step moment estimator, the clever covariate estimator, and the weighted regression estimator. Each augmented weighting method is discussed with a focus on how survey weights are integrated at various stages. Section 3.5 presents the closed-form sandwich variance estimation for the augmented estimators for computationally efficient inference. This section primarily focuses on the development of estimators under the retrospective case, consistent with our previous notation, while modifications required for prospective designs are outlined in the Web Appendix A.6.
@@ 3.1 Augmented weighting estimating equations for survey observational data
To improve efficiency over weighting alone, we extend the asymptotic properties of augmented estimators[21] following §13.5 of Tsiatis [22] to formally incorporate survey weights under the retrospective design. Throughout this section, $\tau_{h}$ denotes the target estimand in (2), $\widehat{\tau}_{h}^{\text{AW}}$ denotes the augmented weighting estimator, and $\phi_{h,i}^{\text{AW}}$ denotes the influence function contribution from subject $i$, so that $\sqrt{n}\{\widehat{\tau}_{h}^{\text{AW}}-\tau_{h}\}=n^{-1/2}\sum_{i=1}^{n}\phi_{h,i}^{\text{AW}}+o_{p}(1)$. In Web Appendix A.4, we first establish the following result.
@@
If the survey weights $1/p_{z}(\mathbf{X})$ and the propensity scores $e_{\text{sp}}(\mathbf{X})$ are known, the class of observed sample data influence functions for regular asymptotically linear estimators of population estimand $\tau_{h}$ is,
$\phi_{h}^{\text{AW}}$ $=\frac{h(\mathbf{X})/p(\mathbf{X})}{\Psi}\left[\frac{ZYp(\mathbf{X})}{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}-\frac{(1-Z)Yp(\mathbf{X})}{p_{0}(\mathbf{X})(1-e_{\text{sp}}(\mathbf{X}))}\right]-\tau_{h}+\left\{Z-\frac{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}{p(\mathbf{X})}\right\}\varphi(\mathbf{X}),$ (10)
for any function $\varphi(\mathbf{X})$, where $\Psi=E[h(\mathbf{X})/p(\mathbf{X})\mid S=1]$.
@@
Here $\Psi$ is the normalization term and the tilting function $h(\mathbf{X})$ can depend on the population propensity score $e_{\text{sp}}(\mathbf{X})$, as illustrated in Table 1. The augmentation term $\{Z-p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})/p(\mathbf{X})\}\varphi(\mathbf{X})$ has mean zero and captures the additional efficiency gain from outcome modeling. Based on this proposition, one can follow the argument in Tsiatis [22] and show that, among all estimators whose influence functions fall within this class, the augmented estimator with
$\varphi(\mathbf{X})$ $=-\frac{h(\mathbf{X})/p(\mathbf{X})}{\Psi}\left[\frac{p(\mathbf{X})E[Y\mid Z=1,\mathbf{X},S=1]}{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}+\frac{p(\mathbf{X})E[Y\mid Z=0,\mathbf{X},S=1]}{p_{0}(\mathbf{X})(1-e_{\text{sp}}(\mathbf{X}))}\right],$ (11)
@@
achieves the smallest asymptotic variance, where $E[Y\mid Z=z,\mathbf{X},S=1]$ is modeled by correctly specified outcome models for $z=1$ and 0. This motivates the augmented weighting estimating equation tailored to survey observational data, where the estimator $\widehat{\tau}_{h}^{\text{AW}}$ solves:
$\frac{1}{\Psi}\frac{1}{n}\sum_{i=1}^{n}\Bigg\{$ $\frac{h(\mathbf{X}_{i})}{p(\mathbf{X}_{i})}(m_{1}(\mathbf{X}_{i})-m_{0}(\mathbf{X}_{i}))+\frac{Z_{i}h(\mathbf{X}_{i})}{e_{\text{sp}}(\mathbf{X}_{i})p_{1}(\mathbf{X}_{i})}(Y_{i}-m_{1}(\mathbf{X}_{i}))$
$-\frac{(1-Z_{i})h(\mathbf{X}_{i})}{(1-e_{\text{sp}}(\mathbf{X}_{i}))p_{0}(\mathbf{X}_{i})}(Y_{i}-m_{0}(\mathbf{X}_{i}))-\tau_{h}\Bigg\}=0,$ (12)
where $m_{z}(\mathbf{X})=E[Y\mid\mathbf{X},Z=z,S=1]$ represents the outcome regression for $z=1$ and 0. By substituting $(\omega_{1,p}(\mathbf{X}),\omega_{0,p}(\mathbf{X}))$ defined in Equation (5) and $\Psi=E[h(\mathbf{X})/p(\mathbf{X})\mid S=1]=E[Z\omega_{1,p}(\mathbf{X})\mid S=1]=E[(1-Z)\omega_{0,p}(\mathbf{X})\mid S=1]$ into above Equation (3.1), the solution to this estimating equation is the augmented weighting estimator,
$\widehat{\tau}_{h}^{\text{AW}}$ $=\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})\left(\widehat{m}_{1}(\mathbf{X}_{i})-\widehat{m}_{0}(\mathbf{X}_{i})\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})}$ (13)
$\quad+\frac{\sum_{i=1}^{n}Z_{i}\omega_{1,p}(\mathbf{X}_{i})\left(Y_{i}-\widehat{m}_{1}(\mathbf{X}_{i})\right)}{\sum_{i=1}^{n}Z_{i}\omega_{1,p}(\mathbf{X}_{i})}-\frac{\sum_{i=1}^{n}(1-Z_{i})\omega_{0,p}(\mathbf{X}_{i})\left(Y_{i}-\widehat{m}_{0}(\mathbf{X}_{i})\right)}{\sum_{i=1}^{n}(1-Z_{i})\omega_{0,p}(\mathbf{X}_{i})}.$
The first term represents the outcome-model-based prediction of $\tau_{h}$, while the second and third terms provide bias correction through weighted residuals when the outcome models are misspecified. Under retrospective designs, when the marginalized survey weights $1/p(\mathbf{X}_{i})$ are not directly accessible, they can be approximated by $p(\mathbf{X}_{i})=r_{z}(\mathbf{X}_{i})p_{z}(\mathbf{X}_{i})$ for each individual with $z_{i}=1$ or 0, as defined in Section 2.1. It is worth noting that estimating $r_{z}(\mathbf{X})$ requires specification of not only the population-level propensity score $e_{\text{sp}}(\mathbf{X})$ but also the sample-level propensity score $e_{\text{fp}}(\mathbf{X})$, whose form becomes a nonlinear logistic under retrospective treatment-dependent sampling when the true population-level treatment assignment follows a standard logistic model, as shown in Equation (1). In Web Appendix B.3, we propose two strategies to resolve this issue: (i) a practical approximation using standard unweighted logistic regression, which can be justified under low sampling rates (generally below 10%) typical of large-scale surveys where log-linear and logistic models are approximately equivalent, and (ii) a nonlinear maximum likelihood approach using the induced form derived from Equation (1), which requires explicit modeling of the sampling mechanism and may be preferred when the sampling rate is much higher.